# A Simple Unified Uncertainty-Guided Framework for Offline-to-Online Reinforcement Learning

### v2

@@ Abstract
Offline reinforcement learning (RL) provides a promising solution to learning an agent fully relying on a data-driven paradigm. However, constrained by the limited quality of the offline dataset, its performance is often sub-optimal. Therefore, it is desired to further finetune the agent via extra online interactions before deployment. Unfortunately, offline-to-online RL can be challenging due to two main challenges: constrained exploratory behavior and state-action distribution shift. To this end, we propose a Simple Unified uNcertainty-Guided (SUNG) framework, which naturally unifies the solution to both challenges with the tool of uncertainty. Specifically, SUNG quantifies uncertainty via a VAE-based state-action visitation density estimator. To facilitate efficient exploration, SUNG presents a practical optimistic exploration strategy to select informative actions with both high value and high uncertainty. Moreover, SUNG develops an adaptive exploitation method by applying conservative offline RL objectives to high-uncertainty samples and standard online RL objectives to low-uncertainty samples to smoothly bridge offline and online stages. SUNG achieves state-of-the-art online finetuning performance when combined with different offline RL methods, across various environments and datasets in D4RL benchmark.  [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
@@
Offline reinforcement learning (RL) Levine et al. (2020), which enables agents to learn from a fixed dataset without interacting with the environment, has demonstrated significant success in many tasks where interaction is expensive or risky Burns et al. (2023); Lee et al. (2023); Zhao et al. (2023). However, such a data-driven learning paradigm inherently limits the agent’s performance as it fully relies on a typically sub-optimal dataset Lee et al. (2022); Zhang et al. (2023). To overcome this issue, the RL community has been exploring the offline-to-online setting, which incorporates additional online interactions to further finetune the pretrained offline RL agents Lee et al. (2022); Mao et al. (2022); Zhang et al. (2023); Zheng et al. (2022).
@@
While the pretraining-finetuning paradigm in offline-to-online RL is natural and intuitive, delivering the expected online improvement can be challenging in practice due to two main challenges: (1) Constrained exploratory behavior: The conservative offline RL objectives, which require the agents to perform actions within the support of the dataset during pretraining, can constrain the agents’ exploratory behavior for online interactions. As a result, they cannot achieve efficient exploration, and thus fail to fully benefit from the trial-and-error paradigm in online RL. (2) State-action distribution shift: During finetuning stage, agents may encounter unfamiliar state-action regimes that fall outside of the support of the dataset. Accordingly, the state-action distribution shift between offline and online data occurs, leading to the well-known extrapolation error Fujimoto et al. (2019); Kumar et al. (2020) during exploitation, which may wipe out the good initialization obtained from the pretraining stage. Existing research typically focuses on tackling one aspect of the challenges above by developing either efficient exploration Zheng et al. (2022); Mark et al. (2022) or effective exploitation Lee et al. (2022); Zheng et al. (2023); Zhang et al. (2023), resulting in limited offline-to-online improvement. However, simultaneously addressing both challenges brings additional difficulty due to the inherent conflict between exploration and exploitation: aggressive exploration may exacerbate the distribution shift, while conservative exploitation can hinder agents from efficient online finetuning.
@@
In this work, we aim to provide a unified solution to both challenges for better sample efficiency during online finetuning. To achieve this, we recognize that both challenges are closely tied to appropriate identification and treatment of unseen state-action pairs. Specifically, efficient exploration necessitates discovery of informative unseen regions of state-action space. Furthermore, effective exploitation needs precise characterization of out-of-distribution (OOD) data to avoid either overly conservative or aggressive policy learning. As such, we leverage proper quantification and utilization of the uncertainty Lee et al. (2021); Bai et al. (2022); An et al. (2021); Wu et al. (2021) to naturally address both challenges and achieve the desired trade-off between exploration and exploitation.
@@
To this end, we present a Simple Unified uNcertainty Guided (SUNG) framework, which provides a generic solution for offline-to-online RL to enable finetuning agents pretrained with different offline RL objectives. Unlike recent uncertainty-aware RL methods that rely on the ensemble technique for uncertainty quantification Bai et al. (2022); An et al. (2021); Lee et al. (2021), SUNG utilizes a simple yet effective approach that estimates the state-action visitation density with a variational auto-encoder (VAE) Kingma and Welling (2013) to quantify the state-action uncertainty. In contrast to prior offline-to-online RL methods, SUNG simultaneously addresses both challenges by leveraging the tool of uncertainty. Concretely, we develop a practical optimistic exploration strategy that follows the principle of optimism in the face of uncertainty Brafman and Tennenholtz (2002); Audibert et al. (2009). The main idea here is to select those state-action pairs with both high value and high uncertainty for efficient exploration. We also propose an adaptive exploitation method to handle the state-action distribution shift by identifying and constraining OOD samples. The key insight is to leverage conservative offline RL objectives for high-uncertainty samples, and standard online RL objectives for low-uncertainty samples. This enables agents to smoothly adapt to changes in the state-action distribution. Moreover, we utilize offline-to-online replay buffer (OORB) Zheng et al. (2023) to incorporate offline and online data during finetuning. Notably, the insights of SUNG are generally applicable and can be combined with most model-free offline RL methods in principle. Empirically, SUNG benefits from consideration of uncertainty to guide both exploration and exploitation, thereby exceeding state-of-the-art methods on D4RL benchmarks Fu et al. (2020), with 14.54% and 14.80% extra averaged offline-to-online improvement compared to the best baseline in MuJoCo and AntMaze domains, respectively.
Summary of Contributions. (1) We propose a generic framework SUNG for sample-efficient offline-to-online RL, which can be combined with existing offline RL methods, such as TD3+BC Fujimoto and Gu (2021) and CQL Kumar et al. (2020). (2) We introduce an optimistic exploration strategy via bi-level action selection to select informative actions for efficient exploration. (3) We develop an adaptive exploitation method with OOD sample identification and regularization to smoothly bridge offline RL and online RL objectives. (4) Experimental results demonstrate that SUNG outperforms the previous state-of-the-art when combined with different offline RL methods, across various types of environments and datasets.
2 Related Work
@@
Offline RL. Offline RL Levine et al. (2020) aims at learning a policy solely from a fixed offline dataset. Most prior works in offline RL have been dedicated to addressing the extrapolation error due to querying the value function with OOD actions Fujimoto et al. (2019); Kumar et al. (2020). As such, it is critical to constrain the learned policy to perform actions within the support set of the offline dataset. Common strategies include policy constraint Fujimoto et al. (2019); Kumar et al. (2019); Fujimoto and Gu (2021); Wu et al. (2022), value regularization Kumar et al. (2020); Lyu et al. (2022); Kostrikov et al. (2021); Bai et al. (2022), etc. However, previous works have shown that the the performance of offline RL agent is typically limited by the quality of the datasets Nair et al. (2020); Lee et al. (2022), which prompts further investigation for the offline-to-online setting.
@@
Offline-to-Online RL. Offline-to-online RL involves pretraining with offline RL and finetuning via online interactions. Some offline RL methods naturally support further online finetuning with continual offline RL objectives Wu et al. (2022); Lyu et al. (2022); Kostrikov et al. (2022). However, they typically tend to be too conservative and yield limited performance improvement Yu and Zhang (2023). Besides, some offline-to-online RL approaches are designed for one specific offline RL method Zheng et al. (2022); Nakamoto et al. (2023); Zhao et al. (2022); Luo et al. (2023). For example, ODT Zheng et al. (2022) introduces the max-entropy RL for finetuning DT Chen et al. (2021). ABCR Zhao et al. (2022) proposes to adaptively loosen the conservative objectives of TD3+BC Fujimoto and Gu (2021). In contrast, we focus on the generic offline-to-online RL framework that can be combined with different offline RL methods Lee et al. (2022); Mark et al. (2022); Zheng et al. (2023); Zhang et al. (2023). Prior works typically address the challenges of exploration limitation and state-action distribution shift. For the former, O3F Mark et al. (2022) utilizes the knowledge in value functions to guide exploration, but it is not compatible with value regularization based offline RL methods. For the latter, BR Lee et al. (2022) utilizes the ensemble technique together with a balanced replay buffer that prioritizes near-on-policy samples. APL Zheng et al. (2023) proposes to take different advantages of offline data and online data for adaptive policy learning. PEX Zhang et al. (2023) freezes the pretrained policies and trains a new policy from scratch using an adaptive exploration strategy to avoid erasing pre-trained policies. Unlike the aforementioned approaches, SUNG unifies the two challenges by emphasizing the proper estimation and utilization of uncertainty.
Uncertainty for RL. Uncertainty-aware RL has achieved notable success in both online and offline RL Lockwood and Si (2022). In terms of online RL, a prominent topic in exploration is the study of epistemic uncertainty, which leads to the development of the well-known principle of optimism in the face of uncertainty Brafman and Tennenholtz (2002); Audibert et al. (2009); Ciosek et al. (2019); Chen et al. (2017); Lee et al. (2021). In the context of offline RL, both model-based Kidambi et al. (2020); Guo et al. (2022); Yu et al. (2020) and model-free An et al. (2021); Bai et al. (2022); Wu et al. (2021); Ghasemipour et al. (2022) approaches can benefit from the consideration of uncertainty to identify and handle OOD actions. However, the aforementioned methods typically utilize the ensemble technique to estimate the uncertainty, which can be computationally expensive in terms of both time and training resources. In contrast, we adopt VAE Kingma and Welling (2013) for efficient and effective uncertainty quantification. Additionally, we investigate the offline-to-online setting, which distinguishes our approach from them.
3 Preliminaries
@@
RL. We follow the standard RL setup that formulates the environment as a Markov decision process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p(s^{\prime}|s,a)$ is the transition distribution, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. The goal of RL is to find a policy $\pi(a|s)$ that maximizes the expected return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$.
@@
Off-policy RL. Off-policy RL methods, such as TD3 Fujimoto et al. (2018) and SAC Haarnoja et al. (2018), have been widely applied due to their sample efficiency. These methods typically alternate between policy evaluation and policy improvement. In particular, given an experience replay dataset $\mathcal{D}$, TD3 learns a deterministic policy $\pi_{\phi}(s)$ and a state-action value function $Q_{\theta}(s,a)$, parameterized by $\phi$ and $\theta$, respectively. The value function can be updated via temporal difference (TD) learning as
@@
Offline RL. In the offline RL setting, the agent only has access to a fixed dataset $\mathcal{D}=\{(s,a,r,s^{\prime})\}$. Although off-policy RL methods can learn from data collected by any policy in principle, they fail in the offline RL setting. This can be attributed to the well-known extrapolation error Fujimoto et al. (2019); Kumar et al. (2020) due to querying the value function with OOD actions. As such, one main line of model-free offline RL research is to constrain the learned policy to perform actions within the support of the dataset in different ways, such as policy constraint Fujimoto and Gu (2021); Wu et al. (2022), value regularization Kumar et al. (2020); Kostrikov et al. (2021); Lyu et al. (2022) and etc. Among them, two representative offline RL methods are TD3+BC Fujimoto and Gu (2021) and CQL Kumar et al. (2020). The former adds a behavior cloning (BC) regularization term to the standard policy improvement in TD3:
@@
where $\lambda_{\rm{BC}}$ balances the standard policy improvement loss and BC regularization. In contrast, the latter resorts to pessimistic under-estimate of Q values during policy evaluation in SAC by
$\mathcal{L}_{Q}^{\rm{CQL}}(\theta)=\mathcal{L}_{Q}(\theta)+\lambda_{\rm{CQL}}\left(\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\phi}}\left[Q_{\theta}\left(s,a\right)\right]+\mathbb{E}_{(s,a)\sim\mathcal{D}}\left[-Q_{\theta}\left(s,a\right)\right]\right),$ (4)
@@
where $\mathcal{F}$ represents either a policy or a value function, parameterized by $\Theta$, $\mathcal{R}$ denotes a regularizer to prevent the extrapolation error, and $\lambda$ balances standard loss and regularization. Note that most model-free offline RL methods can be summarized as Eq.(5) in an explicit Kumar et al. (2020); Fujimoto and Gu (2021); Kostrikov et al. (2021); Wu et al. (2022); Lyu et al. (2022) or implicit manner Kumar et al. (2019); Rezaeifar et al. (2022); Kostrikov et al. (2022); Bai et al. (2022); Xu et al. (2023), which motivates the adaptive exploitation in Section 4.3.
@@
Figure 1: Overview of SUNG. During online finetuning, we alternate between (a) optimistic exploration strategy to collect behavior data from environment and (b) adaptive exploitation method to improve the policy. Moreover, we adopt VAE for state-action density estimation to quantify uncertainty, and OORB to store both offline and online data. This figure is best viewed in color.
4 SUNG

### v3

@@ Abstract
Offline reinforcement learning (RL) provides a promising solution to learning an agent fully relying on a data-driven paradigm. However, constrained by the limited quality of the offline dataset, its performance is often sub-optimal. Therefore, it is desired to further finetune the agent via extra online interactions before deployment. Unfortunately, offline-to-online RL can be challenging due to two main challenges: constrained exploratory behavior and state-action distribution shift. In view of this, we propose a Simple Unified uNcertainty-Guided (SUNG) framework, which naturally unifies the solution to both challenges with the tool of uncertainty. Specifically, SUNG quantifies uncertainty via a VAE-based state-action visitation density estimator. To facilitate efficient exploration, SUNG presents a practical optimistic exploration strategy to select informative actions with both high value and high uncertainty. Moreover, SUNG develops an adaptive exploitation method by applying conservative offline RL objectives to high-uncertainty samples and standard online RL objectives to low-uncertainty samples to smoothly bridge offline and online stages. SUNG achieves state-of-the-art online finetuning performance when combined with different offline RL methods, across various environments and datasets in D4RL benchmark. Codes are made publicly available in https://github.com/guosyjlu/SUNG.
I Introduction
@@
Offline reinforcement learning (RL) [1, 2], which enables agents to learn from a fixed dataset without interacting with the environment, has demonstrated significant success in many tasks where interaction is expensive or risky [3, 4, 5]. However, such a data-driven learning paradigm inherently limits the agent’s performance as it fully relies on a typically sub-optimal dataset [6, 7]. To overcome this problem, the RL community has been exploring the offline-to-online setting, which incorporates additional online interactions to further finetune the pretrained offline RL agents [6, 7, 8].
@@
While the pretraining-finetuning paradigm in offline-to-online RL is natural and intuitive, delivering the expected online improvement can be challenging in practice due to two main challenges: (1) Constrained exploratory behavior: The conservative offline RL objectives, which require the agents to perform actions within the support of the dataset during pretraining, can constrain the agents’ exploratory behavior for online interactions. As a result, they cannot achieve efficient exploration, and thus fail to fully benefit from the trial-and-error paradigm in online RL. (2) State-action distribution shift: During finetuning stage, agents may encounter unfamiliar state-action regimes that fall outside of the support of the dataset. Accordingly, the state-action distribution shift between offline and online data occurs, leading to the well-known extrapolation error [9, 10] during exploitation, which may wipe out the good initialization obtained from the pretraining stage. Existing research typically focuses on tackling one aspect of the challenges above by developing either efficient exploration [8, 11] or effective exploitation [6, 12, 7], resulting in limited offline-to-online improvement. However, simultaneously and efficiently addressing both challenges brings additional difficulty: aggressive exploration may exacerbate the distribution shift, while conservative exploitation can hinder agents from efficient online finetuning.
@@
In this work, we aim to provide a unified solution to both challenges for better sample efficiency during online finetuning. To achieve this, we recognize that both challenges are closely tied to appropriate identification and treatment of unseen state-action pairs. Specifically, efficient exploration necessitates discovery of informative unseen regions of state-action space. Furthermore, effective exploitation needs precise characterization of out-of-distribution (OOD) data to avoid either overly conservative or aggressive policy learning. As such, we leverage proper quantification and utilization of the uncertainty [13, 14, 15, 16] to naturally address both challenges above and achieve the desired trade-off between exploration and exploitation.
@@
To this end, we present a Simple Unified  [^1] Although our framework can be applied to different offline RL algorithms, we still need to manually adjust a few configurations to accommodate different backbone algorithms, which we detail in Section IV-B. uNcertainty Guided (SUNG) framework, which provides a generic solution for offline-to-online RL to enable finetuning agents pretrained with different offline RL objectives. Unlike recent uncertainty-aware RL methods that rely on the ensemble technique for uncertainty quantification [14, 15, 13], SUNG utilizes a simple yet effective approach that estimates the state-action visitation density with a variational auto-encoder (VAE) [17] to quantify the state-action uncertainty. In contrast to prior offline-to-online RL methods, SUNG simultaneously addresses both challenges by leveraging the tool of uncertainty. Concretely, we develop a practical optimistic exploration strategy that follows the principle of optimism in the face of uncertainty [18, 19]. The main idea here is to select those state-action pairs with both high value and high uncertainty for efficient exploration. We also propose an adaptive exploitation method to handle the state-action distribution shift by identifying and constraining OOD samples. The key insight is to leverage conservative offline RL objectives for high-uncertainty samples, and standard online RL objectives for low-uncertainty samples. This enables agents to smoothly adapt to changes in the state-action distribution. Moreover, we utilize offline-to-online replay buffer (OORB) [12] to incorporate offline and online data during finetuning.
Notably, the insights of SUNG are generally applicable and can be combined with most model-free offline RL methods in principle. Empirically, SUNG benefits from consideration of uncertainty to guide both exploration and exploitation, thereby exceeding state-of-the-art methods on D4RL benchmarks [20], with 14.54% and 14.93% extra averaged offline-to-online improvement compared to the best baseline in MuJoCo and AntMaze domains, respectively. In addition, SUNG demonstrates remarkable robustness in finetuning performance across a range of hyper-parameters. Furthermore, we showcase SUNG’s seamless integration with advanced techniques, such as ensemble Q-learning [6, 21] and high update-to-data (UTD) ratio [22], leading to superior finetuning performance.
In summary, our main contributions are four-fold:
  • We propose a generic framework SUNG for sample-efficient offline-to-online RL, flexibly integrating with existing offline RL methods.
  • We introduce an optimistic exploration strategy via bi-level action selection to select informative actions for efficient exploration.
  • We develop an adaptive exploitation method with OOD sample identification and regularization to smoothly bridge offline RL and online RL objectives.
  • Experimental results demonstrate that SUNG outperforms the previous state-of-the-art when combined with different offline RL methods, across various types of environments and datasets.
II Related Work
II-A Offline RL
@@
Offline RL [1] aims at learning a policy solely from a fixed offline dataset. Most prior works in offline RL have been dedicated to addressing the extrapolation error due to querying the value function with OOD actions [9, 10, 23, 24, 25]. As such, it is critical to constrain the learned policy to perform actions within the support set of the offline dataset. Common strategies include policy constraint [9, 26, 27, 28], value regularization [10, 29, 30, 14], etc. However, previous works have shown that the the performance of offline RL agent is limited by the quality of the datasets [31, 6], which motivates investigation for the offline-to-online setting.
II-B Offline-to-Online RL
@@
Offline-to-online RL involves pretraining with offline RL and finetuning via online interactions. Some offline RL methods naturally support further online finetuning with continual offline RL objectives [28, 29, 32]. However, they typically tend to be too conservative and yield limited performance improvement [33]. Besides, some offline-to-online RL approaches are designed for one specific offline RL method [34, 35, 36]. For example, ODT [8] introduces the max-entropy RL for finetuning DT [37]. ABCR [38] proposes to adaptively loosen the conservative objectives of TD3+BC [27]. In contrast, we focus on the generic offline-to-online RL framework that can be combined with different offline RL methods [6, 11, 12, 7, 39, 21]. Prior works typically address the challenges of exploration limitation and state-action distribution shift. For the former, O3F [11] utilizes the knowledge in value functions to guide exploration, but it is not compatible with value regularization based offline RL methods. For the latter, BR [6] utilizes the ensemble technique together with a balanced replay buffer that prioritizes near-on-policy samples. APL [12] proposes to take different advantages of offline data and online data for adaptive policy learning. PEX [7] freezes the pretrained policies and trains a new policy from scratch using an adaptive exploration strategy to avoid erasing pre-trained policies. Unlike them, SUNG unifies both challenges by emphasizing the proper estimation and utilization of uncertainty.
II-C Uncertainty for RL
Uncertainty-aware RL has achieved notable success in reinforcement learning [40], with uncertainty typically categorized as either aleatoric or epistemic. Aleatoric uncertainty arises from the inherent stochasticity of the environment and agent interactions. Distributional RL [41, 42, 43] models this by predicting the return distribution, rather than the exact Q-value. In contrast, epistemic uncertainty originates from the insufficient knowledge and data of the environment, where Bayesian uncertainty quantification methods are widely applied. For example, ensembling based methods [13, 15] utilize the variance of predictions between different models to quantify the uncertainty. Bootstrapped learning based methods [14, 44] extend this idea by training these models on different resampled subsets of the data, further enhancing diversity and robustness. Similarly, Monte-Carlo dropout based methods [45, 16] estimate uncertainty by performing multiple stochastic forward passes through a network with dropout enabled at inference time, and computing the variance across these outputs. Recently, generative models, such as VAE [17, 46] and diffusion models [47, 48], are also applied for uncertainty quantification. Different from the aforementioned works, we focus on the offline-to-online RL setting, and we discuss detailed differences from them in Appendix A in the supplementary material.
III Preliminaries
@@
RL. We follow the standard RL setup that formulates the environment as a Markov decision process (MDP) $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $p(s^{\prime}|s,a)$ is the transition distribution, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. The goal of RL is to find a policy $\pi(a|s)$ that maximizes the expected return $\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})]$. We measure this objective by a state-action value function
$Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}{\gamma^{t}r_{t}}|s_{0}=s,a_{0}=a\right],$ (1)
which calculates the expected discounted return after taking the action $a$ in state $s$.
@@
Off-policy RL. Off-policy RL methods, such as TD3 [49] and SAC [50], have been widely applied due to their sample efficiency. These methods typically alternate between policy evaluation and policy improvement. In particular, given an experience replay dataset $\mathcal{D}$, TD3 learns a deterministic policy $\pi_{\phi}(s)$ and a state-action value function $Q_{\theta}(s,a)$, parameterized by $\phi$ and $\theta$, respectively. The value function can be updated via temporal difference (TD) learning as
@@
Offline RL. In the offline RL setting, the agent only has access to a fixed dataset $\mathcal{D}=\{(s,a,r,s^{\prime})\}$. Although off-policy RL methods can learn from data collected by any policy in principle, they fail in the offline RL setting. This can be attributed to the well-known extrapolation error [9, 10] due to querying the value function with OOD actions. As such, one main line of model-free offline RL research is to constrain the learned policy to perform actions within the support of the dataset in different ways, such as policy constraint [27, 28], value regularization [10, 30, 29] and etc. Among them, two representative offline RL methods are TD3+BC [27] and CQL [10]. The former adds a behavior cloning (BC) regularization term to the standard policy improvement in TD3:
@@
where $\lambda_{\rm{BC}}$ balances the standard policy improvement loss and BC regularization. In contrast, the latter resorts to pessimistic under-estimate of Q values during policy evaluation in SAC:
$\mathcal{L}_{Q}^{\rm{CQL}}(\theta)=\mathcal{L}_{Q}(\theta)$ (5)
$+\lambda_{\rm{CQL}}\left(\mathbb{E}_{s\sim\mathcal{D},a\sim\pi_{\phi}}\left[Q_{\theta}\left(s,a\right)\right]+\mathbb{E}_{(s,a)\sim\mathcal{D}}\left[-Q_{\theta}\left(s,a\right)\right]\right),$
@@
where $\mathcal{F}$ represents either a policy or a value function, parameterized by $\Theta$, $\mathcal{R}$ denotes a regularizer to prevent the extrapolation error, and $\lambda$ balances standard loss and regularization. Note that most model-free offline RL methods can be summarized as Eq. (6) in an explicit [10, 27, 30, 28, 29] or implicit manner [26, 51, 32, 14, 52], which motivates the adaptive exploitation in Section IV-C.
Offline-to-Online RL. While offline RL has found widespread application in constrained interaction scenarios, its efficacy is inherently tied to the quality of the offline dataset. Thus, it becomes imperative to engage in subsequent online finetuning before deploying the model. Formally, the concept of offline-to-online RL delves into the study of a scenario where the agent $\pi_{\phi}$, often accompanied by the Q function $Q_{\theta}$, undergoes initial pretraining using the offline dataset $\mathcal{D}$ through a designated offline RL algorithm. Following this pretraining phase, the agent is further finetuned via online interactions with the environment. This online finetuning process employs off-policy RL algorithms to align the training methodology with that employed during offline pretraining. Additionally, the online replay buffer $\mathcal{B}$ is typically initialized with the offline dataset $\mathcal{D}$ to facilitate effective reuse of experiences.
@@
Figure 1: Overview of SUNG. During online finetuning, we alternate between (a) optimistic exploration strategy to collect behavior data from environment and (b) adaptive exploitation method to improve the policy. We adopt VAE for state-action density estimation to quantify uncertainty.
IV SUNG